# Simulating nationwide coupled disease and fear spread in an agent-based model

### v1

@@ 4.3 Comparing selected scenarios in EpiCast
We observe that fear levels peak after 35 to 60 days at around 60-70% of the population in most cases, with the exception of the pure-fear withdrawal scenario without broadcasters (hosp+sick+fear), where fear levels peak after 85 days at around 38% of the population. In most cases, fear levels fall dramatically from this peak before the end of the simulation. The scenarios with reduced susceptibility due to fear both buck this trend, with fear levels only falling slightly by day 200 without broadcasters (hosp+sick+reduced_sus) and dropping slowly by around 15 percentage points in total with broadcasters (hosp+sick+reduced_sus+bc). We note that, in the absence of broadcasters, the reduced susceptibility case (orange dashed line in Fig. 2(b)) displays the highest fear levels for most of the simulation and the pure-fear withdrawal case the lowest (red dashed line in Fig. 2(b)). In the presence of broadcasters, fear trends more closely resemble other runs. Fear levels fall significantly after their peak in the susceptibility reduction case (blue line in Fig. 2(b)), when they remaining roughly constant in the absence of broadcasters. Conversely, fear levels rose much higher before falling in the pure-fear withdrawal case (green dotted line in Fig. 2(b)), peaking at above 60% of the population rather than less than 40%.
In both scenarios with broadcasters, more broadcasters spread fear than countered it throughout the majority of the run, though in the pure-fear withdrawal case this changed shortly before the end of the run. Far more broadcasters were spreading fear in the susceptibility reduction scenario than in the pure-fear withdrawal scenario for the entire duration (compare blue and green lines in Fig. 2(c)), while the reverse is true for fear countering for most of the duration of the run (Fig. 2(d)). Broadcasters were not used in the other scenarios.
@@ 4.4 EpiCast sensitivity analysis
In order to get a better sense of how varying the behavioral response to fear within EpiCast influences the overall trajectory of disease outbreaks, as per RQ 3, we ran a number of simulation varying two influential parameters. The first was the likelihood that an agent who is afraid of the disease – but not currently showing symptoms – withdraws from their regular schedule of activities purely due to fear, . The second was the scaling factor applied to an afraid agent’s susceptibility to infection, , which represents the impact of afraid agents complying with NPIs that do not directly change their daily schedule, such as masking, maintaining social distance while in public, and hand-washing.
The results of conducting this parameter sweep with only local fear spread are shown in Fig. 4. Fig. 4(a) displays the overall trajectories of each outbreak, with the proportion of new cases shown in orange, that of afraid agents show in blue, and epidemic peaks indicated by the dashed vertical lines. Note that all parameter combinations produce a single epidemic wave. The simulation which comes the closest corresponds to and . Further experiments, shown in Fig. 4(b), revealed that a very narrow range of values – , and – produce multiple waves, but that a deviation by as little in either parameter can be sufficient to return to the single wave scheme.
@@
Higher values of correspond to fewer afraid agents and result reduced variation in fear trajectories for different values of . For , we observe that fear levels rapidly plateau for and rise to a peak before falling to varying degrees for . This drop in fear levels is more pronounced for greater values of , when the overall attack rate is greater.
The results of conducting this parameter sweep with both local and broadcaster-based fear spread are shown in Fig. 5. The overall trajectories of each outbreak are shown in Fig. 5(a) with the dashed vertical lines indicating where peaks occur, and the orange and blue lines indicating proportion of new cases and afraid agents, respectively. Fig. 5(b) highlights which combinations of parameters produce one and two epidemic waves. All single wave scenarios occur with low values of , and either a high or low value of (i.e the upper right corner of Fig. 5(b) and the top two cells in the left of that figure). All other parameter combinations produce two waves. These two waves are generally about the same height with but the second is both taller and longer than the first in most other cases.
@@ 5 Discussion
Examining the fear levels in each scenario, shown in Fig. 2(b), offers some insights into why we might observe multiple peaks primarily in cases with non-local fear spread via broadcasters. In cases where fear reduces agent susceptibility to infection, in the absence of fear spread through broadcasters (hosp+sick+reduced_sus) fear levels never fall far below their peak around 50 days. This is likely due to infection levels never rising to the point where there is a sufficient number of agents in the state to cause afraid agents to lose their fear. This is in contrast to what occurs with fear spread through broadcasters (hosp+sick+reduced_sus+bc), where fear levels quickly fall after reaching their peak. This suggests that the influence of broadcasters is sufficient to lower fear levels to the point where the disease can start spreading again, causing a second wave. Roughly the opposite appears to occur in the cases where we allow purely fear-based withdrawals; without the influence of broadcasters (hosp+sick+fear), fear levels remain lower than in any other scenario for the duration of the run. This is likely because withdrawals reduce both fear and disease spread in in this mode there is a high chance that any given afraid person will withdraw. Conversely, with the influence of broadcasters (hosp+sick+fear+bc) fear levels rise much higher before falling, though they remain lower than in all other scenarios except that with only hospitalization-based withdrawals. This suggests broadcasters allow fear to spread even when most fearful agents remain withdrawn.
@@
These explanations are supported by examining the number of broadcasters spreading (see Fig. 2(c)) and countering fear spread (see Fig. 2(d)) over time. The period from about 25-50 where fear levels peak and then begin to fall in the reduced susceptibility scenario with broadcasters is also the period when we see the number of broadcasters countering fear spread dramatically increase before reaching a much lower sustained growth rate which again corresponds to a slow linear decrease in the number of afraid agents. The number of fear spreading broadcasters follows a similar trend to fear levels, here, albeit with a smaller falloff after the peak. Similarly, we see overall fear levels and the number of fear-spreading broadcasters in the pure-fear withdrawal case, with the dramatic reduction in fearful agents between days 50 and 100 matched by decline in fear-spreading broadcasters of similar magnitude.
@@
In addition, the very narrow range of parameter values for which we do observe multiple waves with purely local fear spread appears to fall in the midst of a phase transition between a regime of small, short outbreaks and large, long outbreaks as represented by the cell , and the cell in Fig. 4(a). The former regime corresponds to lower values of which represent a strong NPI response to fear, implying that fear-induced behavior reduces disease spread sufficiently to forestall a larger outbreak here. The latter regime, by contrast, represents minimal behavioral responses to fear, as seen by the fact it centers on , where fear has no impact of the behavior of agents without symptoms. Intriguingly, these two regimes also appear with broadcaster-based fear spread, and correspond to the only cells that do not exhibit multiple waves in Fig. 5(a).
@@
Altogether, these observations suggest that we observe two waves for opposite reasons in these cases. With reduced susceptibility, fear levels generally do not fall far enough for disease transmission to pick back up and produce a second wave without broadcasters countering fear spread. Conversely, with pure-fear fear-based withdrawals, fear levels generally fail to rise high enough fast enough for the first wave to leave a large reservoir of susceptible agents without broadcasters increasing the rate of fear spread. In either case, non-local fear spread, as represented by broadcasters, appears to facilitate a transition from high initial fear levels to low final fear levels in the presence of fear-based behaviors that significantly reduce disease spread. We posit that this transition only occurs under a few narrow range of conditions in the case of purely local fear spread as a result of fear both spreading and receding relatively slowly. In contrast, the faster rate at which fear levels rise and fall in the presence of broadcasters may explain why multiple epidemic waves are more commonly observed under these conditions. In other words, in order to produce multiple epidemic waves in this model, fear must first spread significantly faster than the disease and then recede faster than it as well. Within this framing, it is clear why broadcasters facilitate the formation of multiple waves: information (and the behavioral changes it engenders) both spreads and is forgotten more rapidly when it is less tightly constrained by physical distance.
@@
The results we present in this paper are perhaps best understood as an exploration of the high-level implications of our model. Though the foundational elements of Epicast are mature and have been calibrated extensively, there remain significant obstacles to doing so for the fear-based components of the model. While some survey data is available describing various behaviors throughout part of the pandemic [27], it is more difficult to tie specific behaviors to specific influences or rationales. We thus confine ourselves to presenting the model itself and exploring the impact of varying parameter values, and leave calibration efforts to future work.
@@
In addition, we acknowledge that the parameter sweep based sensitivity analysis study presented above is necessarily incomplete. While we explore the full possible range of parameter values for and , we do so at a relatively coarse resolution, so as to limit the total number of simulation runs. As evidenced by the multi-wave scenarios presented in Fig. 4(b), this coarseness means that this study may miss dynamics that only emerge for a small range of parameter values. In addition, there are many simulation parameters beyond the two we chose to study. Extending the parameter sweep method to the other model parameters would be prohibitively expensive, but in the absence of such an experiment it is unclear how varying those parameters might affect the dynamics in Epicast.
@@ 6 Conclusion
In this work we presented a series of models which capture the coupled spread of fear of a disease and the disease itself. We do so with several ODE models in order to demonstrate the impact of adding additional disease states to the model, representing asymptomatic, exposed, and pre-symptomatic disease states, and an ABM covering all disease states with both local and non-local mechanisms for fear spread implemented.
@@
In our ABM implementation in EpiCast, we primarily observe multiple epidemic waves when using our non-local fear spread mechanism, which represents traditional broadcast media. We find that a combination of (1) high initial fear levels, (2) a strong behavioral response to fear, and (3) low fear levels later in the simulation produces multiple waves in our model. Thus, one potential explanation for why we only observe multiple waves with purely local fear spread for a very small range of parameter choices is that fear generally spreads (and recedes) much more slowly in the absence of broadcasters in our model.
@@ Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and by cooperative agreement CDC-RFA-FT-23-0069 from the CDC’s Center for Forecasting and Outbreak Analytics. Its contents are solely the responsibility of the authors and do not necessarily represent the official views of the Centers for Disease Control and Prevention.
This work was performed at Los Alamos National Laboratory (LANL), an equal opportunity employer, which is operated by Triad National Security, LLC, for the National Nuclear Security Administration (NNSA) of the U.S. Department of Energy (DOE) under contract #19FED1916814CKC. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. The findings and conclusions in this report are those of the authors and do not necessarily represent the official position of LANL. This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE).
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021
@@
J.K., P.C.A, S.Y.V., and T.C.G. developed the conceptual framework and J.K. wrote the majority of the code for the behavioral models presented in this work. This code included the ODE models, and the extension to EpiCast, with P.C.A. contributing some code to the later. P.C.A. and T.C.G. provided insights into the original design and implementation of EpiCast which informed the implementation of the behavioral model. J.K., A.B., S.Y.V., and T.C.G. conceived the experiments, and J.K. conducted them and analyzed the results. J.K. prepared the initial draft of the manuscript, along with all figures. All authors reviewed the manuscript.
@@
We intend to release all output data from the simulations described in this paper available via Zenodo link upon publication, along with all scripts necessary to generate the plots contained within this manuscript from those results. These scripts are unfortunately still in the process of being reviewed for release by the project sponsors and are thus not available at this time.
@@
Supplementary Material
S1 Coupled contagions ODE models
We provide the full formulation of the system of ODEs which define the model below. This system is summarized in the flow diagram in Fig. S1. Let be the size of the compartment with disease state and fear state . The system is as follows:
@@
Parameter values used in ODE model comparison experiments (Section 4.2) are shown in Table S1 while those used in the EpiCast scenario comparision experiments (Section 4.3) are shown in Table S2.
Variable | Description | Experiment | ||||||||
(a) | (b) | (c) | (d) | (e) | (f) | |||||
disease transmission rate | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | ||||
disease recovery rate | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
0.6 | 0.6 | 0.6 | 0.6 | ||||||
0.4 | 0.4 | 0.4 | 0.4 | ||||||
0.75 | 0.75 | 0.75 | 0.75 | ||||||
0.5 | 0.5 | ||||||||
0.22 | 0.22 | 0.22 | 0.22 | 0.22 | 0.22 | ||||
baseline fear loss rate | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | ||||
fear loss contact rate | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | ||||
1 | 0 | 0 | 0 | 0 | 0 | ||||
1 | 1 | 1 | 1 | 1 | 1 | ||||
0.25 | 0.25 | 0.25 | 0.35 | 0.25 | 0.35 |
Variable | Description | Scenario | ||||||||
(a) | (b) | (c) | (d) | (e) | (f) | |||||
disease transmission rate | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | ||||
disease recovery rate | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
0.6 | 0.6 | 0.6 | 0.6 | 0.6 | 0.6 | ||||
0.4 | 0.4 | 0.4 | 0.4 | 0.4 | 0.4 | ||||
0.75 | 0.75 | 0.75 | 0.75 | 0.75 | 0.75 | ||||
0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
0.22 | 0.22 | 0.22 | 0.22 | 0.22 | 0.22 | ||||
baseline fear loss rate | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | ||||
fear loss contact rate | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | ||||
0 | 0 | 0 | 0 | 0 | 0 | ||||
1 | 1 | 1 | 1 | 1 | 1 | ||||
1 | 1 | 1 | 1 | 0.35 | 0.35 | ||||
0 | 1 | 1 | 1 | 1 | 1 | ||||
0 | 0 | 0.65 | 0.65 | 0 | 0 | ||||
0 | 0 | 0 | 0.25 | 0 | 0.25 | ||||
0.5 | 0.5 | ||||||||
0.25 | 0.25 | ||||||||
0.125 | 0.125 |

### v2

@@ 4.3 Comparing selected scenarios in EpiCast
We observe that fear levels peak after 35 to 60 days at around 60-70% of the population in most cases, with the exception of the pure-fear withdrawal scenario without broadcasters (hosp+sick+fear), where fear levels peak after 85 days at around 38% of the population. In most cases, fear levels fall dramatically from this peak before the end of the simulation. The scenarios with reduced susceptibility due to fear both buck this trend, with fear levels only falling slightly by day 200 without broadcasters (hosp+sick+reduced_sus) and dropping slowly by around 15 percentage points in total with broadcasters (hosp+sick+reduced_sus+bc). We note that, in the absence of broadcasters, the reduced susceptibility case (hosp+sick+reduced_sus in Fig. 2(b)) displays the highest fear levels for most of the simulation and the pure-fear withdrawal case the lowest (hosp+sick+fear in Fig. 2(b)). In the presence of broadcasters, fear trends more closely resemble other runs. Fear levels fall significantly after their peak in the susceptibility reduction case (hosp+sick+reduced_sus+bc in Fig. 2(b)), when they remaining roughly constant in the absence of broadcasters. Conversely, fear levels rose much higher before falling in the pure-fear withdrawal case (hosp+sick+fear+bc in Fig. 2(b)), peaking at above 60% of the population rather than less than 40%.
In both scenarios with broadcasters, more broadcasters spread fear than countered it throughout the majority of the run, though in the pure-fear withdrawal case this changed shortly before the end of the run. Far more broadcasters were spreading fear in the susceptibility reduction scenario than in the pure-fear withdrawal scenario for the entire duration (compare hosp+sick+reduced_sus+bc and hosp+sick+fear+bc in Fig. 2(c)), while the reverse is true for fear countering for most of the duration of the run (Fig. 2(d)). Broadcasters were not used in the other scenarios.
Figure 2: EpiCast Fear Spread Scenarios: New cases (2(a)), total fear levels (2(b)), total broadcasters spreading fear (2(c)), and total broadcasters countering fear spread (2(d)) for six different scenarios run in EpiCast. These scenarios can include agents withdrawing from their normal schedules due to being hospitalized (hosp), being fearful of the disease while having symptoms of it (sick), and fear of the disease without symptoms (fear), as well as non-local fear spread through broadcast media (bc), and fearful agents having a lower susceptibility to the disease due to taking protective actions such as masking (reduced_sus).
@@
Figure 3: Geographic Distribution of Outbreaks in Two Epicast Fear Spread Scenarios: Geographic distribution of new case counts for selected timestamps (left) and overall new case trends for selected states (right) for scenarios with pure-fear withdrawals without (upper) and with (lower) broadcaster-based fear spread.
@@ 4.4 EpiCast sensitivity analysis
In order to get a better sense of how varying the behavioral response to fear within EpiCast influences the overall trajectory of disease outbreaks, as per RQ 3, we ran a number of simulation varying two influential parameters. The first was the likelihood that an agent who is fearful of the disease – but not currently showing symptoms – withdraws from their regular schedule of activities purely due to fear, . The second was the scaling factor applied to an fearful agent’s susceptibility to infection, , which represents the impact of fearful agents adopting protective behaviors that do not directly change their daily schedule, such as masking, maintaining social distance while in public, and hand-washing.
The results of conducting this parameter sweep with only local fear spread are shown in Fig. 4. Fig. 4(a) displays the overall trajectories of each outbreak, with the proportion of new cases shown in orange, that of fearful agents show in blue, and epidemic peaks indicated by the dashed vertical lines. Note that all parameter combinations produce a single epidemic wave. The simulation which comes the closest corresponds to and . Further experiments, shown in Fig. 4(b), revealed that a very narrow range of values – , and – produce multiple waves, but that a deviation by as little in either parameter can be sufficient to return to the single wave scheme.
@@
Higher values of correspond to fewer fearful agents and result reduced variation in fear trajectories for different values of . For , we observe that fear levels rapidly plateau for and rise to a peak before falling to varying degrees for . This drop in fear levels is more pronounced for larger values of , when the overall attack rate is higher.
Figure 4: Epicast local fear spread sensitivity analysis: New cases by day (with peaks indicated by dashed vertical line) for (4(a)) a coarse-grained parameter search and (4(b)) a finer-grained search highlighting cases with multiple waves, and (4(c)) total attack rate for various rates of purely fear-based withdrawals () and levels of susceptibility scaling due to fear () with only local fear spread.
The results of conducting this parameter sweep with both local and broadcaster-based fear spread are shown in Fig. 5. The overall trajectories of each outbreak are shown in Fig. 5(a) with the dashed vertical lines indicating where peaks occur, and the orange and blue lines indicating proportion of new cases and fearful agents, respectively. Fig. 5(b) highlights which combinations of parameters produce one and two epidemic waves. All single wave scenarios occur with low values of , and either a high or low value of (i.e the upper right corner of Fig. 5(b) and the top two cells in the left of that figure). All other parameter combinations produce two waves. These two waves are generally about the same height with but the second is both taller and longer than the first in most other cases.
@@
Figure 5: Epicast broadcaster fear spread sensitivity analysis: (5(a)) new cases by day (with peaks indicated by dashed vertical line), (5(b)) number of epidemic waves, and (5(c)) total attack rate for various rates of purely fear-based withdrawals () and levels of susceptibility scaling due to fear () with both local and broadcaster-based fear spread.
@@ 5 Discussion
Examining the fear levels in each scenario, shown in Fig. 2(b), offers some insights into why we might observe multiple peaks primarily in cases with non-local fear spread via broadcasters. In cases where fear reduces agent susceptibility to infection, in the absence of fear spread through broadcasters (hosp+sick+reduced_sus) fear levels never fall far below their peak, at around 50 days. This is likely due to infection levels never rising to the point where there is a sufficient number of agents in the state to cause fearful agents to lose their fear. This is in contrast to what occurs with fear spread through broadcasters (hosp+sick+reduced_sus+bc), where fear levels quickly fall after reaching their peak. This suggests that the influence of broadcasters is sufficient to lower fear levels to the point where the disease can start spreading again, causing a second wave. Roughly the opposite appears to occur in the cases where we allow purely fear-based withdrawals; without the influence of broadcasters (hosp+sick+fear), fear levels remain lower than in any other scenario for the duration of the run. This is likely because withdrawals reduce both fear and disease spread in this mode and there is a high chance that any given fearful person will withdraw. Conversely, with the influence of broadcasters (hosp+sick+fear+bc) fear levels rise much higher before falling, though they remain lower than in all other scenarios except that with only hospitalization-based withdrawals. This suggests broadcasters allow fear to spread even when most fearful agents remain withdrawn.
@@
These explanations are supported by examining the number of broadcasters spreading (see Fig. 2(c)) and countering fear spread (see Fig. 2(d)) over time. The period from days 25-50 where fear levels peak and then begin to fall in the reduced susceptibility scenario with broadcasters is also the period when we see the number of broadcasters countering fear spread dramatically increase, before reaching a much slower sustained growth rate. This slower growth period corresponds to a slow linear decrease in the number of fearful agents. The number of fear spreading broadcasters follows a similar trend to fear levels in this case, albeit with a smaller falloff after the peak. Similarly, we see that overall fear levels and the number of fear-spreading broadcasters show similar patterns in the pure-fear withdrawal case, with the dramatic reduction in fearful agents between days 50 and 100 matched by decline in fear-spreading broadcasters of similar magnitude.
@@
In addition, the very narrow range of parameter values for which we do observe multiple waves with purely local fear spread appears to fall in the midst of a phase transition between a regime of small, short outbreaks and large, long outbreaks as represented by the cell , and the cell in Fig. 4(a), respectively. The former regime corresponds to lower values of which represent a strong NPI response to fear, implying that fear-induced behavior reduces disease spread sufficiently to forestall a larger outbreak here. The latter regime, by contrast, represents minimal behavioral responses to fear, as seen by the fact it centers on , where fear has no impact of the behavior of agents without symptoms. Intriguingly, these two regimes also appear with broadcaster-based fear spread, and correspond to the only cells that do not exhibit multiple waves in Fig. 5(a).
@@
Altogether, these observations suggest that we observe two waves for opposite reasons in these cases. With reduced susceptibility, fear levels generally do not fall far enough for disease transmission to pick back up and produce a second wave without broadcasters countering fear spread. Conversely, with pure-fear fear-based withdrawals, fear levels generally fail to rise high enough fast enough for the first wave to leave a large reservoir of susceptible agents without broadcasters increasing the rate of fear spread. In either case, non-local fear spread, as represented by broadcasters, appears to facilitate a transition from high initial fear levels to low final fear levels in the presence of fear-based behaviors that significantly reduce disease spread. We posit that this transition only occurs under a few narrow range of conditions in the case of purely local fear spread as a result of fear both spreading and receding relatively slowly. In contrast, the faster rate at which fear levels rise and fall in the presence of broadcasters may explain why multiple epidemic waves are more commonly observed under these conditions. In other words, in order to produce multiple epidemic waves in this model, fear must first spread significantly faster than the disease and then recede faster than it as well. Within this framing, it is clear why broadcasters facilitate the formation of multiple waves: information and the behavioral changes it engenders both spreads and recedes more rapidly when it is less tightly constrained by physical distance.
@@
The results we present in this paper are perhaps best understood as an exploration of the high-level implications of our model. Though the foundational elements of EpiCast are mature and have been calibrated extensively, there remain significant obstacles to doing so for the fear-based components of the model. While some survey data is available describing various behaviors throughout part of the pandemic [29], it is much more difficult to tie behavior to specific influences or rationales. We thus confine ourselves to presenting the model itself and exploring the impact of varying parameter values, and leave calibration efforts to future work.
@@
In addition, we acknowledge that the parameter sweep based sensitivity analysis study presented above is necessarily incomplete. While we explore the full possible range of parameter values for and , we do so at a relatively coarse resolution, so as to limit the total number of simulation runs. As evidenced by the multi-wave scenarios presented in Fig. 4(b), this coarseness means that this study may miss dynamics that only emerge for a small range of parameter values. In addition, there are many simulation parameters beyond the two we chose to study. Extending the parameter sweep method to the other model parameters would be prohibitively computationally expensive, but in the absence of such an experiment it is unclear how varying those parameters might affect the dynamics in EpiCast.
@@ 6 Conclusion
In this work we present a series of models which capture the coupled spread of fear of a disease and the disease itself. We do so with several ODE models in order to demonstrate the impact of adding additional disease states to the model, representing asymptomatic, exposed, and pre-symptomatic disease states, and an ABM covering all disease states with both local and non-local mechanisms for fear spread implemented.
@@
In our ABM implementation in EpiCast, we primarily observe multiple epidemic waves when using our non-local fear spread mechanism, which represents broadcast media (regional information communication, including print, radio, and TV). We find that a combination of (1) high initial fear levels, (2) a strong behavioral response to fear, and (3) low fear levels later in the simulation produce multiple waves in our model. Thus, one potential explanation for why we only observe multiple waves with purely local fear spread for a very small range of parameter choices is that fear generally spreads (and recedes) much more slowly in the absence of broadcasters in our model.
@@ Acknowledgments
This work was supported by the U.S. Department of Energy,
Office of Science, Office of Advanced Scientific Computing
Research and by cooperative agreement CDC-RFA-FT-23-0069 from
the CDC’s Center for Forecasting and Outbreak Analytics.
Its contents are solely the responsibility of the authors and
do not necessarily represent the official views of the Centers
for Disease Control and Prevention.
This work was performed at
Los Alamos National Laboratory (LANL), an equal opportunity
employer, which is operated by Triad National Security, LLC,
for the National Nuclear Security Administration
(NNSA) of the U.S. Department of Energy (DOE) under contract
#19FED1916814CKC. The funders had no role in study design,
data collection and analysis, decision to publish, or preparation
of the manuscript. The findings and conclusions in this report
are those of the authors and do not necessarily represent
the official position of LANL.
This research used resources provided by the Darwin testbed at
LANL which is funded by the Computational Systems and Software
Environments subprogram of LANL’s Advanced Simulation and
Computing program (NNSA/DOE).
This material is based upon work supported by the U.S. Department
of Energy, Office of Science, Office of Advanced Scientific
Computing Research, Department of Energy Computational Science
Graduate Fellowship under Award Number DE-SC0021.
@@
J.K., P.C.A, D.J.B., C.B., S.Y.D.V., and T.C.G. developed the conceptual framework and J.K. wrote the majority of the code for the behavioral models presented in this work. This code included the ODE models, and the extension to EpiCast, with P.C.A. contributing some code to the later. D.J.B. and C.B. provided suggestions regarding the development and evaluation of the ODE models. P.C.A. and T.C.G. provided insights into the original design and implementation of EpiCast which informed the implementation of the behavioral model. J.T. developed and provided access to UrbanPop, the synthetic population on which this study was conducted, and guided that use. J.K., A.B., S.Y.D.V., and T.C.G. conceived the experiments, and J.K. conducted them and analyzed the results. J.K. prepared the initial draft of the manuscript, along with all figures. A.B., S.Y.D.V., and T.C.G. all provided supervision. All authors reviewed and edited the manuscript.
@@
We have publicly released all output data from simulations described in this paper through Zenodo here: https://zenodo.org/uploads/15706577. The source code for the scripts used to analyze these results and produce the plots shown in this manuscript are publicly available here: https://github.com/lanl/epicast-coupled. While we hope to eventually release the full code of both EpiCast and the fear spread extension presented in this manuscript, this process is especially long for patentable code given the policies of Los Alamos National Laboratory (where EpiCast is developed) and the multiple U.S. government agencies which sponsored this work. Limited access to this code may be granted upon request.
@@
Parameter values used in ODE model comparison experiments (Section 4.2) are shown in Suplementary Table S1 while those used in the EpiCast scenario comparision experiments (Section 4.3) are shown in Suplementary Table S2.
Variable
Description
Experiment
disease transmission rate
0.2
0.2
0.2
0.2
0.2
0.2
disease recovery rate
0.5
0.5
0.5
0.5
0.5
0.5
probability of
symptomatic infection
0.6
0.6
0.6
0.6
probability of
asymptomatic infection
0.4
0.4
0.4
0.4
relative infectivity of
asymptomatic infection
0.75
0.75
0.75
0.75
inverse of
incubation period
0.5
0.5
fear transmission
rate
0.22
0.22
0.22
0.22
0.22
0.22
baseline fear loss rate
0.05
0.05
0.05
0.05
0.05
0.05
fear loss contact rate
0.44
0.44
0.44
0.44
0.44
0.44
relative fear
susceptibility after
symptomatic recovery
1
0
0
0
0
0
relative infectivity
when fearful
1
1
1
1
1
1
relative susceptibility
when fearful
0.25
0.25
0.25
0.35
0.25
0.35
Table S1: Parameter values used in ODE model experiments: Results of experiments (a-e) are shown in the corresponding subfigures in Fig. 1, and the EpiCast scenarios to setups with (a) only hospitalization-based withdrawals, (b) the addition of symptomatic fearful withdrawals, and the addition of either pure-fear withdrawals from fear – without (c) or with (d) broadcaster-based fear spread – or reduced susceptibility when fearful – without (e) or with (f) broadcaster-based fear spread.
Variable
Description
Scenario
disease transmission rate
0.2
0.2
0.2
0.2
0.2
0.2
disease recovery rate
0.5
0.5
0.5
0.5
0.5
0.5
probability of
symptomatic infection
0.6
0.6
0.6
0.6
0.6
0.6
probability of
asymptomatic infection
0.4
0.4
0.4
0.4
0.4
0.4
relative infectivity of
asymptomatic infection
0.75
0.75
0.75
0.75
0.75
0.75
inverse of
incubation period
0.5
0.5
0.5
0.5
0.5
0.5
fear transmission
rate
0.22
0.22
0.22
0.22
0.22
0.22
baseline fear loss rate
0.05
0.05
0.05
0.05
0.05
0.05
fear loss contact rate
0.44
0.44
0.44
0.44
0.44
0.44
relative fear
susceptibility after
symptomatic recovery
0
0
0
0
0
0
relative infectivity
when fearful
1
1
1
1
1
1
relative susceptibility
when fearful
1
1
1
1
0.35
0.35
probability of withdrawal
when fearful with symptoms
0
1
1
1
1
1
probability of pure-fear
withdrawal when fearful
0
0
0.65
0.65
0
0
probability of watching
a given broadcaster
0
0
0
0.25
0
0.25
threshold of fearful
workers for broadcasters
to take position
0.5
0.5
relative rate of new cases
for broadcasters to
resume neutral position
0.25
0.25
relative rate of new cases
for broadcasters
to counter fear spread
0.125
0.125
Table S2: Parameter values used in EpiCast scenario experiments: Parameters represent scenarios with (a) only hospitalization-based withdrawals, (b) the addition of symptomatic fearful withdrawals, and the addition of either pure-fear withdrawals – without (c) or with (d) broadcaster-based fear spread – or reduced susceptibility when fearful – without (e) or with (f) broadcaster-based fear spread.